OpenStudy (anonymous):

The equation of the line of symmetry pf the curve y=x^(2)+px+21 is x=4. Find the value of p.

OpenStudy (anonymous):

is p=-8

OpenStudy (anonymous):

stpes?

OpenStudy (anonymous):

here is y-5=(x-4)^2

OpenStudy (anonymous):

for x=4 the line of symmetry thr rhs of the equation should have (x-4)^2 =x^2-2(4)x+16 hence coefficient of x is -8 thus p=-8

OpenStudy (anonymous):

what is it ? how come? y-5=(x-4)^2

OpenStudy (anonymous):

hey...